OpenStudy (anonymous):

How would I find the inverse of a question like f(x)=10x+4...?

OpenStudy (valpey):

\[f(x)=10x+4 = y\] \[10x = y-4\] \[x = \frac{y-4}{10} = g(y)\] g is the inverse function of f